OpenStudy (anonymous):

where did i go wrong o_O For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8 -2y + 6 2y - 6 2y + 6 -2y - 6

OpenStudy (anonymous):

i multipied top equation by 3 to elimanate X then left 7y = -8

OpenStudy (jdoe0001):

well... your elimination is.... close, no there, but close but you're asked to do it by using substitution and not to solve it, just to get the expression that will be used in the 2nd by substituting the expression from the 1st one

OpenStudy (anonymous):

ahh so -2y one

OpenStudy (anonymous):

but -6 or + 6 so we want to cpmpletely wipe it out? if so add 6

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} -x + 2y = -6&\implies 2y+6={\color{brown}{ x}}\\ \hline\\ 3{\color{brown}{ x}} + y = 8&\implies 3({\color{brown}{ 2y+6}})+y=8 \end{array}\)

OpenStudy (solomonzelman):

\[-x + 2y = -6\] 1) subtract 2y from both sides (to isolate variable x) 2) divide both sides by -1 (to make it x=....)

OpenStudy (anonymous):

so would it be a ) -2y + 6

OpenStudy (solomonzelman):

No

OpenStudy (anonymous):

im confused then ?

OpenStudy (anonymous):

wats r=the answer?

OpenStudy (jdoe0001):

\(\bf -x + 2y = -6\qquad {\color{brown}{ +x}} \\ \quad \\ \cancel{ {\color{brown}{ +x}}-x } + 2y = -6{\color{brown}{ +x}}\qquad {\color{brown}{ -6}} \\ \quad \\ {\color{brown}{ -6}}+2y=\cancel{ -{\color{brown}{ -6}}+6 }+x\)

OpenStudy (jdoe0001):

you'd solve for "x" on the 1st equation and substitute THAT X on the 2nd equation\(\large \begin{array}{llll} -x + 2y = -6&\implies 2y+6={\color{brown}{ x}}\\ \hline\\ 3{\color{brown}{ x}} + y = 8&\implies 3({\color{brown}{ 2y+6}})+y=8 \end{array}\)

OpenStudy (anonymous):

im confused here

OpenStudy (jdoe0001):

?

OpenStudy (jdoe0001):

ok....what part?

OpenStudy (anonymous):

we add -2 to cancel the ys

OpenStudy (anonymous):

but dpnt get wat next

OpenStudy (jdoe0001):

you might be thinking about the "elimination" procedure, but you're asked to use "substitution" instead

OpenStudy (anonymous):

are we trying to get rid of x or y ?

OpenStudy (anonymous):

can u tell me answe and ill try work it out

OpenStudy (jdoe0001):

\(\bf \text{For the following system,}\\ \text{ if you isolated x in the first equation to use }\\ {\color{red}{ \text{ the Substitution Method }}} , \\ \textit{what expression would you substitute into the second equation?}\)

OpenStudy (anonymous):

i thought u could on sub in a solved wquation

OpenStudy (jdoe0001):

well... you could, yes, since they all correlate, it's just that the exercise only requires you to do it in certain way

OpenStudy (anonymous):

so a is the right answer?

OpenStudy (anonymous):

plz mayun , so tired

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

what expression would you substitute into the second equation? <--- meaning, what would you get for "x" from the 1st equation

OpenStudy (jdoe0001):

\(\bf -x + 2y = -6\qquad {\color{brown}{ +x}} \\ \quad \\ \cancel{ {\color{brown}{ +x}}-x } + 2y = -6{\color{brown}{ +x}}\qquad {\color{brown}{ -6}} \\ \quad \\ {\color{brown}{ -6}}+2y=\cancel{ -{\color{brown}{ -6}}+6 }+x\)

OpenStudy (anonymous):

o_O really have no idea

OpenStudy (anonymous):

-6 + 2y = x so we need to add 6 andd - 2?

OpenStudy (jdoe0001):

hmm I have an extra -.... but anyhow \(\bf \bf -x + 2y = -6\qquad \qquad {\color{brown}{ +x}} \\ \quad \\ \cancel{ {\color{brown}{ +x}}-x } + 2y = -6{\color{brown}{ +x}}\qquad \qquad {\color{brown}{ -6}} \\ \quad \\ {\color{brown}{ -6}}+2y=\cancel{ {\color{brown}{ -6}}+6 }+x\)

OpenStudy (anonymous):

y is y not byitself?

OpenStudy (anonymous):

thought we isolate y

OpenStudy (anonymous):

X = -6 + 2y

OpenStudy (jdoe0001):

\(\bf \text{For the following system,}\\ \text{ if you isolated } {\huge x} \textit{ in the first equation to use }\\ {\color{red}{ \text{ the Substitution Method }}} , \\ \textit{what expression would you substitute into the second equation?}\)

OpenStudy (jdoe0001):

yeap, x = -6+2y or 2y-6